# Multi-Scale Attention and Gated Shifting for Fine-Grained Event Spotting in Videos

### v1

@@ Abstract
Precise Event Spotting (PES) in sports videos requires frame-level recognition of fine-grained actions from single-camera footage. Existing PES models typically incorporate lightweight temporal modules such as Gate Shift Module (GSM) or Gate Shift Fuse (GSF) to enrich 2D CNN feature extractors with temporal context. However, these modules are limited in both temporal receptive field and spatial adaptability. We propose a Multi-Scale Attention Gate Shift Module (MSAGSM) that enhances GSM with multi-scale temporal dilations and multi-head spatial attention, enabling efficient modeling of both short- and long-term dependencies while focusing on salient regions. MSAGSM is a lightweight plug-and-play module that can be easily integrated with various 2D backbones. To further advance the field, we introduce the Table Tennis Australia (TTA) dataset—the first PES benchmark for table tennis—containing over 4,800 precisely annotated events. Extensive experiments across five PES benchmarks demonstrate that MSAGSM consistently improves performance with minimal overhead, setting new state-of-the-art results.
@@ 1 Introduction
Detecting the precise moment when events occur in a video is a core challenge in video understanding, particularly in the single-camera setting commonly used in most sports recordings. Precise Event Spotting (PES) aims to localize and classify events at frame-level accuracy, typically within a 1–2 frame tolerance [11]. This fine-grained precision is critical in sports scenarios, where even a few frames of deviation can correspond to a different event entirely (see Figure 1). Accurate PES provides a strong foundation for downstream tasks such as ball tracking, player positioning, and automatic highlight generation. Moreover, precise spotting plays a vital role in post-game analysis by surfacing key moments and reducing the need to manually browse full-length videos. For instance, according to Table Tennis Australia’s sports analytics team, annotating events in a single match video can take 4–5 hours. Automating this process not only streamlines analysis but also enables timely feedback for athletes and coaches.
Figure 1: Example of Precise Event Spotting: in table tennis, the moment a player contacts the ball during a serve (red) or when a ball bounces on the table (blue).
@@
Recent advancements in video understanding have explored various approaches to this task, including transformer-based models [1, 41], multi-modal methods [30, 34], and self-supervised frameworks [4]. Among these, E2E-Spot [11] has emerged as a strong baseline, inspiring several subsequent models such as the current state-of-the-art T-DEED [35] and UGLF [29]. A key component in E2E-Spot is the Gate Shift Module (GSM) [25], which enables 2D CNNs to learn temporal dynamics by shifting features across adjacent frames. While effective, GSM is limited in its temporal range and does not explicitly model longer-term dependencies.
@@
Despite these advancements, PES remains a highly challenging task due to its unique characteristics. Events are often difficult to distinguish from adjacent frames—sometimes even for human annotators—without access to extended temporal context [35]. Moreover, different events may require varying amounts of temporal information to be correctly identified, further complicating model design. In addition to temporal reasoning, effective PES also demands accurate spatial attention: in sports videos, a large portion of each frame typically consists of background, making it essential for models to focus on small but informative regions—such as the ball or a player’s hand position—to make correct predictions. Another practical limitation lies in the availability of suitable datasets. Most publicly available sports video datasets are derived from broadcast footage of elite sports [11, 10, 8, 38], which are often professionally edited and filmed from ideal angles. However, in many real-world sports settings—such as local or developmental leagues—matches are recorded using handheld devices with varied perspectives, inconsistent quality, and frequent occlusions.
To address these challenges, we propose the Multi-Scaled Attention Gate Shift Module (MSAGSM), a lightweight module that enhances the temporal modeling capacity of 2D CNNs. MSAGSM builds upon the GSM framework by incorporating multi-scale temporal shifts and attention mechanisms. This design allows the model to capture longer-range dependencies while simultaneously attending to important spatial regions such as the ball or player positions. Importantly, MSAGSM is computationally efficient and can be easily integrated into existing 2D backbones without the resource overhead associated with 3D models.
@@
In addition, we introduce a new dataset for PES: the Table Tennis Australia (TTA) dataset, recorded and curated in collaboration with Paralympics Australia’s sports analysts. The dataset includes 39 games and a total of 4,878 annotated events across eight event types: serve, bounce, forehand, and backhand—each distinguished between the near and far sides of the table. To our knowledge, TTA is the first PES dataset specifically designed for table tennis. Compared to existing PES datasets, TTA is both sizable and challenging. Its scale is comparable to FineGym [22] and surpasses that of FineDiving [38] and FigureSkating [10]. Moreover, unlike tennis [11], table tennis involves more densely packed, fine-grained events within shorter time spans, presenting a more demanding benchmark for precise spotting.
@@
Our main contributions are summarized as follows: (1) We propose MSAGSM, a lightweight module that combines multi-scale temporal shifts with spatial attention, and can be efficiently plugged into existing 2D CNN backbones. (2) We introduce the Table Tennis Australia (TTA) dataset, the first PES dataset for table tennis, containing 39 games and 4,878 annotated events across 8 classes. (3) MSAGSM, when integrated with standard 2D backbones, achieves state-of-the-art performance across multiple PES benchmarks, including TTA.
@@
Figure 2: Overview of our proposed architecture for Precise Event Spotting (PES). The input video is processed through a 2D CNN backbone for spatial feature extraction, followed by the Multi-Scaled Attention Gate Shift Module (MSAGSM) for spatio-temporal modeling. MSAGSM combines channel-wise attention and multi-scale temporal shifting to capture varying temporal dependencies. The resulting features are passed through a GRU-based temporal model to predict event class probabilities at the frame level.
@@ 2 Related Work
Spatial-Temporal Modelling Significant progress has been made in video understanding tasks—such as video classification and action recognition—driven by the development of deep convolutional neural networks (CNNs) and Transformer-based architectures [5, 6, 17, 2, 27]. One major challenge in these models is their high memory and computational requirements, which make them less suitable for deployment on resource-constrained devices compared to lightweight 2D CNNs.
@@
To address this limitation, the Gate Shift Module (GSM) [25] was introduced. Built upon TSM, GSM employs an additional spatio-temporal 3D convolutional kernel as a gating mechanism to selectively decide which features to shift forward or backward based on contextual information. Its variant, Gate Shift Fuse (GSF) [26], adds a point-wise convolution block before and after the gating operation to enable interaction between spatial and spatio-temporal feature groups, further enhancing modeling capability.
@@
Precise Event Spotting In video understanding tasks such as video classification and action recognition, a common approach is to sparsely sample frames from the entire video [7, 31, 13]. In contrast, Precise Event Spotting (PES) requires dense frame sampling—particularly for sports videos—due to the fast motion and high frequency of events occurring within short time spans.
Compared to Temporal Action Detection (TAD), which focuses on predicting temporal intervals of actions [16, 24], PES aims to identify actions using a single key frame. This is especially relevant in sports scenarios, where the start and end boundaries of actions (e.g., a forehand shot) can be visually ambiguous and difficult to define precisely [37].
@@
To address this ambiguity, the SoccerNet dataset introduced the task of Action Spotting (AS) [8], which focuses on identifying key frames of important events within a relaxed temporal window (e.g., $\pm$50 frames). Subsequently, PES was proposed to enforce a stricter temporal tolerance, requiring models to be accurate within just a few frames.
@@
E2E-Spot [11] serves as a strong PES baseline, combining RegNetY with the Gate Shift Module (GSM) for temporal modeling. Building on this, T-DEED [35] addressed the challenge of distinguishing between adjacent frames with highly similar visual appearance. It incorporated the Scalable Granularity Perception (SGP) module [23] to enhance discriminability among temporally close features.
UGLF [29] takes a different direction by incorporating the vision-language model GLIP [14]. GLIP processes both an image and a textual prompt to extract meaningful object-level features such as ’ball’, ’referee’, or ’card’, marking the first use of vision-language models in the PES domain.
@@
Beyond visual features, audio signals can also serve as important cues in sports—for instance, crowd noise after a goal or racket contact sound in tennis. ASTRA [34], a transformer-based encoder-decoder model designed for soccer, integrates both visual and audio embeddings, using learnable queries in the decoder to jointly attend to multi-modal cues.
@@
PES Datasets Existing datasets are often acquired from broadcast videos from high-level matches such as [8, 22, 38, 11, 10]. These datasets are fine-grained, where all event and class labels relate to a single activity.
Despite progress, existing models often fail to capture both short- and long-range temporal dependencies while attending to relevant spatial regions. They also rely on broadcast-quality footage, limiting applicability in grassroots or para-sports where handheld cameras are used. To address this, we propose the Multi-Scaled Attention Gate Shift Module (MSAGSM), a lightweight temporal modeling block, and introduce Table Tennis Australia (TTA)—the first PES benchmark for table tennis, captured in real-world conditions using single-camera recordings by sports analysts.
@@ 3 Multi-Scale Attention GSM
3.1 Overview
Our approach enables 2D CNNs to capture varying spatio-temporal dependencies through an efficient combination of multi-scale gated shifting and attention mechanisms. The proposed method consists of two main components: (1) A multi-head attention module, which splits the input features into multiple groups and applies attention independently to each group, allowing the model to focus on spatially informative regions; (2) A multi-scale GSM, which introduces temporal dilations to shift features across varying temporal distances. This design allows the model to directly access a broader temporal context compared to the original GSM. An overview of the module architecture is illustrated in Figure 2.
3.2 Attention GSM
Our intuition is rooted in the observation that, in sports videos, most critical events are concentrated around key entities such as players, referees, and the ball. In contrast, a large portion of each frame consists of background content that contributes little to event detection. Motivated by the effectiveness of attention mechanisms in prior work [33, 5, 3, 32], we extend the original GSM into a multi-head attention GSM. This design allows the model to focus on spatially informative regions while suppressing irrelevant background, thereby improving the accuracy of event spotting. Figure 3 shows the details of spatial attention,
@@ 3.2 Attention GSM
We generate spatial attention maps using 2D convolutional layers to highlight the most informative regions within a feature map. Given an intermediate feature tensor, we first split it into multiple chunks along the channel dimension, where the number of chunks corresponds to a predefined number of attention heads—a tunable hyperparameter.
For each head, a 2D convolutional layer generates a spatial attention map based on the original features. These maps are applied independently to their corresponding feature chunks, allowing each head to specialize in distinct spatial regions in a data-driven manner. Finally, the attended feature chunks are concatenated to reconstruct a unified, refined feature representation.
3.3 MultiScale GSM
@@
We adopt the GSM as a reference architecture, as it has been proven effective in the PES domain [11, 35, 29]. GSM performs a gating-based temporal shifting mechanism inspired by TSM [15] and GST [19]. However, a key limitation of GSM is that it only performs shifts across adjacent frames, restricting its ability to capture longer-range temporal dependencies directly, which is often important in sports videos to classify events and discriminate similar frames.
@@
We propose the MSAGSM, which extends GSM by replacing fixed, adjacent-frame shifting with a learnable, spatially-aware gating mechanism that enables direct access to longer temporal ranges. The architectural design of MSAGSM is illustrated in Figure 4.
@@
Figure 4: Structure of the proposed Multi-Scale Gate Shift Module (MSGSM). Given input features, a 3D CNN generates gate maps corresponding to different temporal shift ranges (e.g., $\Delta t=\pm 1$, ±$\pm 2$). Features are shifted left and right across multiple scales, gated accordingly, and combined through concatenation and stacking. This design enables the model to adaptively capture short- and long-term temporal dependencies in a data-driven manner.
@@
Given a sequence of feature maps $X\in\mathbb{R}^{C\times T\times H\times W}$, where $C$ is the number of channels, and $T$, $H$, $W$ denote the temporal length, height, and width, respectively. For each temporal dilation scale, MSGSM first applies a gating layer implemented as a 3D CNN. This gating layer receives a temporal window of three feature maps sampled across time (with dilation) centered around the current timestamp. Two gating maps are then generated for each feature map—corresponding to left and right shifts. These gates are activated using the tanh function to constrain their values between (-1, +1), ensuring stable modulation. $(G_{\text{left}},G_{\text{right}})=\tanh(\text{Conv3D}(X))$.
@@
The feature map is split into two groups $X=[X_{\text{left}},X_{\text{right}}]$, each modulated by one of the gates and then temporally shifted left or right via a rolling operation
$\hat{X}_{\text{left}}$ $=\text{Roll}(X_{\text{left}}\odot G_{\text{left}},-\delta)$ (1)
$\hat{X}_{\text{right}}$ $=\text{Roll}(X_{\text{right}}\odot G_{\text{right}},+\delta)$ (2)
@@
Padding is applied to maintain shape integrity and prevent feature loss at the sequence boundaries. To preserve important information, the shifted features are added back to the original input via a residual connection. The gated, shifted features are then concatenated along the channel dimension to restore the original shape. $\hat{X}=[\hat{X}_{\text{left}},\hat{X}_{\text{right}}]+X$
@@
Finally, the outputs from different dilation levels are stacked along a new dimension and passed through a learnable weighting mechanism. This layer adaptively determines the importance of each temporal scale, enabling the model to incorporate information across varying temporal distances. Formally, let $\{\hat{X}^{(i)}\}_{i=1}^{N}$ denote the outputs from $N$ different shift ranges. The final output is computed as:
$\hat{X}_{\text{final}}=\sum_{i=1}^{N}w_{i}\cdot\hat{X}^{(i)},\quad\text{where }\sum_{i=1}^{N}w_{i}=1,\;w_{i}\geq 0$ (3)

### v2

# MSAGSM: Multi-Scale Attention Gated Shifting for Precise Event Spotting

@@ Abstract
Precise Event Spotting (PES) in sports videos requires frame-level recognition of fine-grained actions from single-camera footage. Existing PES models typically incorporate lightweight temporal modules such as the Gate Shift Module (GSM) or the Gate Shift Fuse to enrich 2D CNN feature extractors with temporal context. However, these modules are limited in both temporal receptive field and spatial adaptability. We propose a Multi-Scale Attention Gate Shift Module (MSAGSM) that enhances GSM with multi-scale temporal shifts and channel grouped spatial attention, enabling efficient modeling of both short and long-term dependencies while focusing on salient regions. MSAGSM is a lightweight, plug-and-play module that integrates seamlessly with diverse 2D backbones. To further advance the field, we introduce the Table Tennis Australia dataset, the first PES benchmark for table tennis containing over 4,800 precisely annotated events. Extensive experiments across four PES benchmarks demonstrate that MSAGSM consistently improves performance with minimal overhead, setting new state-of-the-art results. All code and datasets are publicly available at https://anonymous.4open.science/r/MSAGSM-B0A4.
@@ 1 Introduction
Detecting the precise moment when events occur in a video is a core challenge in video understanding, particularly in the single camera setting commonly used in most sports recordings. Precise Event Spotting (PES) aims to localize and classify events at frame-level accuracy, typically within a 1–2 frame tolerance [12]. This fine-grained precision is critical in sports scenarios, where even a few frames of deviation can correspond to a different event entirely (see Figure 1). Accurate PES provides a strong foundation for downstream tasks such as ball tracking, player positioning, and automatic highlight generation. Moreover, precise spotting plays a vital role in post-game analysis by highlighting key moments and reducing the need to manually browse full-length videos. For instance, according to Table Tennis Australia’s sports analytics team, annotating events in a single match video can take 4–5 hours. Automating this process not only streamlines analysis but also enables timely feedback for athletes and coaches.
@@
Recent advancements in video understanding have explored various approaches to this task, including transformer-based models [1, 54], multimodal methods [38, 43], and self-supervised frameworks [5]. Among these, E2E-Spot [12] has emerged as a strong baseline, inspiring several subsequent models such as the current state-of-the-art (SOTA) T-DEED [44], UGLF [37], and ASTRM [27]. A key component in E2E-Spot is the Gate Shift Module (GSM) [31], which enables 2D CNNs to learn temporal dynamics by shifting features across adjacent frames. While effective, GSM is limited in its temporal range and does not explicitly model longer-term dependencies.
@@
Despite these advancements, PES remains a highly challenging task due to its unique characteristics. Events are often difficult to distinguish from adjacent frames, sometimes even for human annotators without access to extended temporal context [44]. Moreover, different events may require varying amounts of temporal information to be correctly identified, further complicating model design. In addition to temporal reasoning, effective PES also demands accurate spatial attention: in sports videos, a large portion of each frame typically consists of background, making it essential for models to focus on small but informative regions such as the ball or a player’s hand position to make correct predictions. Another practical limitation lies in the availability of suitable datasets. Most publicly available sports video datasets are derived from broadcast footage of elite sports [12, 11, 9, 48], which are often professionally edited and filmed from ideal angles. However, in many real-world sports settings, such as local or developmental leagues, matches are recorded using handheld devices with varied perspectives, inconsistent quality, and frequent occlusions.
Figure 1: Example of Precise Event Spotting in table tennis: the exact moments when the player contacts the ball during a serve (highlighted in red) and when the ball bounces on the table (highlighted in yellow). The ball locations are circled accordingly. t represents the timestamp at the frame level.
To address these challenges, we propose the Multi-Scale Attention Gate Shift Module (MSAGSM), a lightweight extension of GSM that overcomes its limitation of accessing only adjacent frames by incorporating multi-scale temporal shifts. To mitigate the noise introduced by larger temporal windows, we further design a channel-grouped spatial attention mechanism that learns spatial masks within channel groups to guide shifting, differing from transformer multi-head attention by avoiding costly query–key–value interactions. This design enables MSAGSM to capture longer-range dependencies while focusing on salient regions (e.g., ball or player positions), and it can be seamlessly integrated into 2D CNN backbones with minimal overhead compared to 3D models.
@@
In addition, we introduce a new dataset for PES: the Table Tennis Australia (TTA) dataset, recorded and curated in collaboration with Paralympics Australia’s sports analysts, providing more realistic recordings compared to existing datasets. The dataset includes 39 games and a total of 4,878 annotated events across eight event types: serve, bounce, forehand, and backhand—each distinguished between the near and far sides of the table. To our knowledge, TTA is the first PES dataset specifically designed for table tennis. Compared to existing PES datasets, TTA is both sizable and challenging. Its scale is comparable to FineGym [28] and surpasses that of FigureSkating [11]. Moreover, unlike tennis [12], table tennis involves more densely packed, fine-grained events within shorter time spans, presenting a more demanding benchmark for PES.
@@
Our main contributions are summarized as follows: (1) proposing MSAGSM, a lightweight module that combines multi-scale temporal shifts with spatial attention, and can be efficiently plugged into existing 2D CNN backbones; (2) introduce the TTA dataset, the first PES dataset for table tennis, containing 39 games and 4,878 annotated events across eight classes; and (3) show that MSAGSM, when integrated with standard 2D backbones, achieves SOTA performance across multiple PES benchmarks, including TTA.
@@
Figure 2: Overview of our proposed architecture for Precise Event Spotting. The input video is processed through a 2D CNN backbone for spatial feature extraction, followed by the Multi-Scale Attention Gate Shift Module (MSAGSM) for spatio-temporal modeling. MSAGSM combines channel-group spatial attention and multi-scale temporal shifting to capture varying temporal dependencies. The resulting features are passed through a GRU-based temporal model to predict event class probabilities at the frame level.
@@ 2 Related Work
Spatial-Temporal Modelling: Significant progress has been made in video understanding tasks, such as video classification and action recognition, driven by the development of deep CNNs and Transformer-based architectures [6, 7, 20, 2, 35]. One major challenge in these models is their high memory and computational requirements, which make them less suitable for deployment on resource-constrained devices compared to lightweight 2D CNNs.
@@
To address this limitation, the GSM [31] was introduced. Built upon TSM, GSM employs an additional spatio-temporal 3D convolutional kernel as a gating mechanism to selectively decide which features to shift forward or backward based on contextual information. Its variant, Gate Shift Fuse (GSF) [32], adds a point-wise convolution block before and after the gating operation to enable interaction between spatial and spatio-temporal feature groups, further enhancing modeling capability.
@@
Precise Event Spotting: In video understanding tasks such as video classification and action recognition, a common approach is to sparsely sample frames from the entire video [8, 39, 15]. In contrast, PES requires dense frame sampling, particularly for sports videos due to the fast motion and high frequency of events occurring within short time spans. Compared to Temporal Action Detection, which focuses on predicting temporal intervals of actions [18, 30], PES aims to identify actions using a single key frame. This is especially relevant in sports scenarios, where the start and end boundaries of actions (e.g., a forehand shot) can be visually ambiguous and difficult to define precisely [46].
@@
To address this ambiguity, the SoccerNet dataset introduced the task of Action Spotting [9], which focuses on identifying key frames of important events within a relaxed temporal window (e.g., $\pm$50 frames). Subsequently, PES was proposed to enforce a stricter temporal tolerance, requiring models to be accurate within just a few frames.
@@
E2E-Spot [12] serves as a strong PES baseline, combining RegNetY [25] with the Gate Shift Module (GSM) [31] for temporal modeling. Building on this, T-DEED [44] addressed the challenge of distinguishing between adjacent frames with highly similar visual appearance by incorporating the Scalable Granularity Perception module [29] to enhance discriminability among temporally close features. UGLF [37] takes a different direction by integrating the vision-language model GLIP [16], marking the first use of vision-language features for PES. Most recently, ASTRM [27] further advanced the field by improving long-range temporal dependency through spatial–temporal feature injection into the feature extractor and proposing the Soft Instance Contrastive (SoftIC) loss to mitigate class imbalance.
@@
Beyond visual features, audio signals can also serve as important cues in sports, for instance, crowd noise after a goal. ASTRA [43], a transformer-based encoder-decoder model designed for soccer, integrates both visual and audio embeddings, using learnable queries in the decoder to jointly attend to multi-modal cues.
Other Sports Video Tasks: Ball tracking methods such as the TrackNet family [13, 33, 4, 26, 47, 19] focus on recovering ball trajectories, with TrackNetV4 introducing motion-aware attention maps from frame differences. While related, tracking localizes objects whereas PES targets the timing and semantics of fine-grained events. These approaches are thus complementary, and accurate trajectories could provide useful cues for future event detection.
@@
Sports Video Datasets: Existing PES datasets are often acquired from broadcast videos from high-level matches such as [9, 28, 48, 12, 11]. These datasets are fine-grained, where all event and class labels relate to a single activity. Other sports-related video datasets include ball tracking [13, 33, 34, 47] and sports video captioning [42, 23, 49, 51].
Despite progress, PES methods still struggle to capture long-range temporal dependencies while attending to fine-grained spatial cues. In addition, most datasets rely on broadcast footage with fixed cameras and ideal conditions, which overlook the challenges of community-level sports. To bridge these gaps, we propose MSAGSM, a lightweight block for enhanced temporal reasoning and spatial focus, and introduce TTA, the first table tennis PES benchmark with consumer-grade courtside videos featuring variable viewpoints, camera motion, and frequent occlusions.
@@ 3 Multi-Scale Attention GSM
3.1 Model Overview
Figure 2 illustrates an overview of our proposed MSAGSM architecture. It enables 2D CNNs to capture varying spatio-temporal dependencies through an efficient combination of multi-scale gated shifting and channel-group spatial attention mechanisms. The proposed method consists of two main components: (1) a multi-head channel-grouped spatial attention module, which splits the input features into multiple groups and applies attention independently to each group, allowing the model to focus on spatially informative regions; and (2) a multi-scale GSM, which introduces multiscale temporal dilations to shift features across varying temporal distances. This design allows the model to directly access a broader temporal context compared to the original GSM.
3.2 Channel Grouped Spatial Attention
Our intuition is rooted in the observation that, in sports videos, most critical events are concentrated around key entities such as players, referees, and the ball. In contrast, a large portion of each frame consists of background content that contributes little to event detection. Motivated by the effectiveness of attention mechanisms in prior work [41, 6, 3, 40], we extend the original GSM into a channel-grouped spatial attention GSM. This design allows the model to focus on spatially informative regions while suppressing irrelevant background, thereby improving the accuracy of event spotting. Figure 3 shows the details of the channel-grouped spatial attention mechanism.
We generate spatial attention maps using 2D convolutional layers to highlight the most informative regions within a feature map. Given an intermediate feature tensor, we first split it into multiple groups along the channel dimension, where the number of groups corresponds to a predefined number of attention heads—a tunable hyperparameter.
For each head, a 2D convolutional layer generates a spatial attention map based on the original features. These maps are applied independently to their corresponding feature chunks, allowing each head to specialize in distinct spatial regions in a data-driven manner. Finally, the attended feature chunks are concatenated to reconstruct a unified refined feature representation.
@@ 3.2 Channel Grouped Spatial Attention
3.3 Multi-Scale GSM
@@
We adopt the GSM as a reference architecture, as it has been proven effective in the PES domain [12, 44, 37]. GSM performs a gating-based temporal shifting mechanism inspired by TSM [17] and GST [22]. However, a key limitation of GSM is that it only performs shifts across adjacent frames, restricting its ability to capture longer range temporal dependencies directly, which is often important in sports videos to classify events and discriminate similar frames.
@@
We propose the MSAGSM, which extends GSM by replacing fixed, adjacent frame shifting with a learnable spatially aware gating mechanism that enables direct access to longer temporal ranges. The architectural design of MSAGSM is illustrated in Figure 4.
@@
Figure 4: Structure of the proposed Multi-Scale Gate Shift Module. Given input features, a 3D CNN generates gate maps corresponding to different temporal shift ranges (e.g., $\Delta t=\pm 1$, $\pm 4$). Features are shifted left and right across multiple scales, gated accordingly, and combined through concatenation and stacking. This design enables the model to adaptively capture both short and long-term temporal dependencies in a data-driven manner, while preserving the original information.
@@
Given a sequence of feature maps $X\in\mathbb{R}^{C\times T\times H\times W}$, where $C$ is the number of channels, $T$ is the temporal length, and $H$ and $W$ denote the height and width of a frame, respectively. For each temporal distance scale, MSAGSM first applies a gating layer implemented as a 3D CNN. This gating layer receives a temporal window of three feature maps sampled across time centered around the current timestamp. Two gating maps, corresponding to the left and right shifts, are generated for each feature map. These gates are passed through the tanh activation function to constrain their values to the range (1,+1)(-1,+1), ensuring stable modulation:
$(G_{\text{left}},G_{\text{right}})=\tanh(\text{Conv3D}(X)).$ (1)
@@
The feature map is split into two groups, $X=[X_{\text{left}},X_{\text{right}}]$, each multiplied element-wise by its corresponding gate map and then temporally shifted left or right via a rolling operation:
$\hat{X}_{\text{left}}$ $=\text{Roll}(X_{\text{left}}\odot G_{\text{left}},-\delta),$ (2)
$\hat{X}_{\text{right}}$ $=\text{Roll}(X_{\text{right}}\odot G_{\text{right}},+\delta).$ (3)
Here, $\odot$ denotes element-wise multiplication, and $\delta$ is the temporal shift distance (in frames).
@@
Padding is applied to maintain shape integrity and prevent feature loss at the sequence boundaries. To preserve important information, the shifted features are added back to the original input via a residual connection. The gated, shifted features are then concatenated along the channel dimension to restore the original shape,
$\hat{X}=[\hat{X}_{\text{left}},\hat{X}_{\text{right}}]+X.$ (4)
@@
Finally, the outputs from different distance levels are stacked along a new dimension and passed through a learnable weighting mechanism. This layer adaptively determines the importance of each temporal scale, enabling the model to incorporate information across varying temporal distances. Formally, let $\{\hat{X}^{(i)}\}_{i=1}^{N}$ denote the outputs from $N$ different shift ranges. The final output is computed as:
$\hat{X}_{\text{final}}=\sum_{i=1}^{N}w_{i}\cdot\hat{X}^{(i)},\quad\text{where }\sum_{i=1}^{N}w_{i}=1,\;w_{i}\geq 0.$ (5)